# CrackSCF: Lightweight Cascaded Fusion Network for Robust and Efficient Structural Crack Segmentation

### v4

@@ Abstract
Accurately segmenting structural cracks at the pixel level remains a major hurdle, as existing methods fail to integrate local textures with pixel dependencies, often leading to fragmented and incomplete predictions. Moreover, their high parameter counts and substantial computational demands hinder practical deployment on resource-constrained edge devices. To address these challenges, we propose CrackSCF, a Lightweight Cascaded Fusion Crack Segmentation Network designed to achieve robust crack segmentation with exceptional computational efficiency. We design a lightweight convolutional block (LRDS) to replace all standard convolutions. This approach efficiently captures local patterns while operating with a minimal computational footprint. For a holistic perception of crack structures, a lightweight Long-range Dependency Extractor (LDE) captures global dependencies. These are then intelligently unified with local patterns by our Staircase Cascaded Fusion Module (SCFM), ensuring the final segmentation maps are both seamless in continuity and rich in fine-grained detail. To comprehensively evaluate our method, we created the challenging TUT benchmark dataset and evaluated it alongside five other public datasets. The experimental results show that the CrackSCF method consistently outperforms the existing methods, and it demonstrates greater robustness in dealing with complex background noise. On the TUT dataset, CrackSCF achieved 0.8382 on F1 score and 0.8473 on mIoU, and it only required 4.79M parameters. The code and dataset is available at https://github.com/Karl1109/CrackSCF.
{IEEEkeywords}
Crack segmentation, lightweight networks, cascaded fusion, local patterns, long-range dependencies
@@ 1 Introduction
Cracks are common defects in industrial structures such as pavements and buildings, which can lead to serious safety issues. Thus, automated detection on edge control devices is crucial for ensuring production safety [1]. Early automated detection methods used traditional image processing techniques for crack extraction [2]. While these methods are easy to implement, their performance in complex scenarios still needs improvement.
Figure 1: (a) shows the performance of our method compared to other methods. (b) shows the different scenarios included in TUT dataset.
@@
In recent years, benefiting from the strong local inductive capability of Convolutional Neural Networks (CNN), CNN-based networks such as Crack500 [3] and SC-OSDA [4] have shown efficient feature extraction capabilities in structural crack segmentation tasks. The spatial invariance of CNNs allows them to detect cracks anywhere in an image by learning similar features at different locations. In labeled images, crack labels are long, narrow structures of white pixels, while the background is large black areas. This challenges the model to distinguish between thin cracks and the background. Nevertheless, networks based on CNNs face challenges in capturing long-range dependencies between crack pixels, leading to segmentation discontinuities.
@@
The success of Vision Transformer (ViT) [5] has demonstrated the advantages of Transformer, with their strong sequence processing capabilities, in handling visual tasks [6]. Unlike CNNs, transformers can capture dependencies between any pixels when processing images, making them more advantageous in handling complex topologies. To harness the advantages of both CNNs and transformers effectively, researchers have proposed networks such as DTrCNet [7] and Crackmer [7], which integrate local and long-range information through U-shaped or residual connections.
Despite the aforementioned methods combining CNN and Transformer have improved performance in pixel-level crack segmentation tasks, they still have some limitations, as follows:
  1. 1.
@@
    The different fusion strategies significantly affect the segmentation performance of CNN-Transformer networks. For instance, UCTNet [8] and CTCS [9] use channel concatenation to fuse local patterns and long-range dependences information. However, concatenation merely stacks the feature information together without deeper feature interaction, failing to understand their correlation and complementarity, resulting in poor segmentation performance on fine cracks and susceptibility to background noise. Additionally, DTrCNet [7] and Crackmer [10] use channel attention to fuse information, which can enhance focus on certain features but neglects spatial information and pixel details in the feature maps, impacting the model’s overall performance.
  2. 2.
@@
    Most networks, in pursuit of performance, overlook the parameter and computational costs brought by numerous convolution operations, this makes them challenging to deploy on resource-constrained control devices. As shown in Fig. 1(a), networks like SFIAN [11], Crackmer [10], and CTCS[9] have high computational parameter counts, indicating that they are difficult to implement on resource-limited devices like smartphones and drones, significantly diminishing their practicality.
@@
To address these issues, we propose a staircase cascaded fusion network that generates high-quality segmentation maps with low computational resources while effectively suppressing background noise. Specifically, to reduce the number of parameters and computational burden, we introduce a lightweight convolution block called LRDS. This module utilizes the principle of low-rank approximation to effectively reduce the computational resources required by convolutional layers, and captures spatial structures efficiently in the low-rank space through depthwise convolution, maintaining good feature perception while requiring very low computational resources. We also propose a lightweight Long-range Dependency Extractor (LDE), which uses a lightweight deformable attention mechanism to effectively capture irregular long-range dependencies between crack pixels with very low computational demands. Additionally, to better leverage the complementarity and correlation of feature maps while ensuring that critical detail information is not missed, we introduce the Staircase Cascaded Fusion Module (SCFM). This module generates feature maps layer by layer and effectively fuses them using a pixel attention mechanism that can finely capture local and long-range information and a channel concatenation mechanism, comprehensively combining local texture information and pixel dependencies.
@@
To comprehensively evaluate the structural crack segmentation performance of the model in various complex environments, we also collected a dataset named TUT, which contains 1408 RGB images. Existing public datasets like DeepCrack [12] and Crack500 [3] include only 1-2 types of scenarios and have relatively simple crack shapes, which are insufficient for thoroughly testing the network’s performance. As shown in Fig. 1(b), the TUT dataset includes a broader range of real structural crack image scenarios, covering plastic runways, bricks, tiles, cement, bitumen, generator blades, metal materials, and underground pipelines. With such a diverse set of image scenarios, this dataset effectively evaluates the network’s generalization and robustness. Furthermore, the images feature occlusions, highly complex backgrounds, and uneven lighting, which further enhances the datasets diversity.
In conclusion, our main contributions are as follows:
  1. 1.
    We proposed a novel CrackSCF network for structural crack segmentation. By introducing a lightweight convolution block, a lightweight long-range dependency extractor, and a staircase cascaded fusion module, this network effectively distinguishes between background and crack pixels with low computational resources, resulting in high-quality segmentation maps.
  2. 2.
    We collected the TUT dataset, which includes structural cracks in eight different scenarios, providing a more comprehensive evaluation of the network’s performance.
  3. 3.
    We carried out comprehensive experiments on five public datasets and the TUT benchmark dataset, demonstrating that our proposed CrackSCF network outperforms existing SOTA methods.
The organization of the subsequent sections of this paper is as follows: In Section 2, existing crack detection models and lightweight networks are reviewed. Section 3 details the components of the proposed network. Section 4 delves into the experimental details and results. Finally, Section 5 provides a summary of the entire paper and highlights future research directions.
2 Related Work
2.1 Crack Segmentation Networks
Recently, a growing number of researchers have utilized CNNs to develop crack segmentation networks, achieving advanced performance. Zhu et al. [13] proposed the BSNBM network, which integrates bidirectional self-rectifying modules with Bayesian inference to enhance crack detection robustness and accuracy under uncertainty. Cheng et al. [11] introduced the Selective Feature Fusion and Irregular-Aware Network (SFIAN) to effectively model irregular crack objects, resulting in clearer backgrounds on various crack images. Lyu et al. [14] proposed a CNN-based crack quantification framework that utilizes voxel-based reconstruction and Bayesian data fusion to achieve accurate crack width estimation.
Benefit from the attention mechanism, transformers are particularly well-suited for handling targets with long-range dependencies and complex global structures, especially for cracks, which often have extensive spans and intricate shapes. Therefore, many researchers are amalgamating the outputs of CNNs and transformers for crack segmentation. Qi et al. [15] developed an end-to-end model based on ViT and level set theory for segmenting defects in bridge pavement, achieving precise segmentation by amalgamating the outputs of two parallel decoders. Feng et al. [16] developed a shale crack segmentation network that combines CNN and Swin Transformer, utilizing Fast Fourier Transform (FFT) to merge feature maps from both, effectively suppressing high-frequency noise in the imaging process.
Although CNN and CNN-Transformer methods have shown promise in crack segmentation, CNNs struggle to capture global context due to their local receptive fields, leading to discontinuous predictions and vulnerability to background noise. CNN-Transformer hybrids improve global perception but often adopt single-branch fusion, limiting their ability to integrate local and long-range cues. Moreover, both lack selective interaction and noise suppression, and their high computational cost hinders deployment on edge devices.
2.2 Lightweight Networks
Deploying deep learning models on edge devices like smartphones and drones is trending. However, these devices have limited computing resources, making it challenging to use models with high parameter counts. Thus, compact model design for lightweight modifications in size and speed is crucial. For example, Howard et al. [17] proposed MobileNet, an efficient model using depthwise separable convolutions to reduce computational load and parameter count, suitable for resource-constrained devices. Ma et al. [18] introduced ShuffleNet, providing practical design principles for efficient network architecture. Additionally, low-rank approximation compression, effective in fields like numerical linear algebra and image processing, is now used to compress deep learning network architectures. Swaminathan et al. [19] introduced SLR, a sparse low-rank method that improves compression rates with sparse SVD matrices.
In recent years, there has been a growing trend among researchers to integrate model lightweighting techniques with crack segmentation models, aiming to facilitate real-time crack detection on edge computing devices. For example, Zhu et al. [20] designed RHACrackNet, a lightweight crack detection network integrating hybrid attention and residual blocks, enabling real-time detection on edge devices. Huang et al. [21] proposed a lightweight fusion attention module for road crack detection, ensuring high-precision crack extraction with rapid detection. Although lightweight models reduce computational cost, their aggressive parameter reduction limits model expressiveness, impairing the capture of intricate patterns and leading to suboptimal segmentation. This trade-off between performance and efficiency undermines their reliability in real-world applications.
@@ 3.1 Network Architecture
As shown in Fig. 2, the CrackSCF network has three main components: a lightweight Multi-scale Feature Extractor (MFE), a lightweight Long-range Dependency Extractor (LDE), and a Staircase Cascaded Fusion Module (SCFM). The overall architecture of the network is shaped like a staircase, with feature maps of different scales progressively fused and upsampled. In this process, the advantages of CNNs and Transformers in feature extraction complement each other.
Specifically, for a given dataset, the images are first input into the MFE, which is divided into four stages. In each stage, the extracted feature maps undergo a Feature Enhancement stage. Then, the LDE processes these enhanced feature maps to obtain pixel sequences rich in long-range dependencies. Subsequently, the SCFM receives these sequences along with the four layers of feature maps extracted by the MFE. After processing through four stages, the resolution is doubled at each stage while the number of channels is halved, ultimately resulting in the segmentation result.
Importantly, as shown in Fig. 3, using the original convolution operations results in high FLOPs and Params, especially in the SCFM module, reaching 73.36G and 9.71M, respectively. Replacing all convolutions with LRDS blocks significantly reduced FLOPs and Params, demonstrating the effectiveness of LRDS in model lightweighting.
Figure 3: Comparison of each module when using our LRDS or original convolution separately. (a) shows the comparison of FLOPs, (b) shows the comparison of parameters.
@@ 3.2 Lightweight Multi-scale Feature Extractor
As shown in Fig. 2, our designed LRDS block adopts a bottleneck-like structure. It comprises a Reduce layer, a depthwise convolution layer, a pointwise convolution layer, and an Expand layer. The input feature map first passes through the Reduce layer, reducing the input channels to $f_{mid}$. It then goes through the depthwise convolution layer and the pointwise convolution layer, keeping the number of channels unchanged. Finally, the Expand layer increases the number of channels to $f_{out}$, producing the final output feature map.
@@
Specifically, in the convolutional layer, we assume that its filter has a spatial size of $k$ and a number of input channels of $c$. In computing the convolutional response, the filter is applied to a sub-tensor of the input tensor of shape $k\times k\times c$. Suppose this sub-tensor is $t\in\mathbb{R}^{k^{2}c+1}$, where we append a bias term at the end. The response at a position in the output, $y\in\mathbb{R}^{e}$, can be expressed as follows:
@@
where $e$ represents the number of filters and $W$ is a $e\times(k^{2}c+1)$ matrix. Assuming the vector $y$ lies in a low-rank subspace, we can write it as $y=U(y-y_{1})+y_{1}$, where $U$ is a $e\times e$ matrix of rank $e_{0}$, with $e_{0}<e$. $y_{1}$ is the mean vector of the response. Therefore, we can write the response as
@@
where $b=y_{1}-Uy_{1}$ is the new bias. The rank of $U$ is $e_{0}$, and it can be decomposed into two $e\times e_{0}$ matrices $J$ and $K$ such that $U=JK^{\top}$. We use $W_{1}=K^{\top}W$ to denote a $e_{0}\times(k^{2}c+1)$ matrix, which corresponds to a set of $e_{0}$ filters. Thus, we can rewrite the response as follows:
@@
In order to efficiently capture the spatial information of feature maps in low-rank space for finer perception of detailed regions, we process the features in LRDS by embedding deep convolution and pointwise convolution in low-rank space. We constructed MFE by replacing all convolution operations with LRDS convolution blocks.
@@ 3.3 Lightweight Long-range Dependency Extractor
$\centering\begin{split}&MSDeformAttn(z_{q},\hat{p}_{q},\{x_{c}\}_{c=1}^{C})=\\ &\sum_{h=1}^{H}W_{h}\Bigg[\sum_{c=1}^{C}\sum_{t=1}^{T}A_{hcqt}\cdot W^{\prime}_{h}x_{c}(\phi_{c}(\hat{p}_{q})+\Delta p_{hcqt})\Bigg]\end{split}\@add@centering$ (4)
@@
As shown in Fig. 4(a), Linear operations account for 90% of the FLOPs and 67% of the Params in the LDE module. In order to reduce the computational resources required by the LDE, we propose a linear transformation method (LR Linear) based on low-rank approximation. In traditional linear layers, the weight matrix connecting the input and output is typically a high-rank matrix, representing a complete connection between input and output channels. By applying low-rank approximation, we decompose the weight matrix into two smaller matrices, reducing the number of parameters.
@@
we perform this transformation through two linear layers. First, we apply a linear transformation $B^{T}$ to the input $x\in\mathbb{R}^{r_{\text{in}}}$, resulting in an intermediate representation $z\in\mathbb{R}^{r}$:
@@
Then, we apply another linear transformation $A$ to $z$, resulting in the output $y\in\mathbb{R}^{r_{\text{out}}}$:
@@
As $r\ll r_{\text{out}},r_{\text{in}}$, the low-rank approximation reduces the original matrix to two smaller matrices, significantly reducing the computational and storage requirements. This leads to a reduction in both the computational complexity and the number of parameters, thus improving training and inference efficiency. As shown in Fig. 4(b), replacing all Linear operations in the LDE with LR Linear reduced the parameters and computational load from 7.18G and 1.05M to 0.14G and 0.02M.
The structure of LDE is shown in Fig. 2. LDE facilitates long-range dependency modeling through the interaction of multi-scale features between the encoder and decoder. Initially, the encoder inputs a multi-scale feature set $\{X_{c}\}_{c=1}^{C}$ (where $C$ represents the number of feature levels). After undergoing layer flattening and positional encoding enhancement, the encoder extracts the feature sequence $S\in\mathbb{R}^{B\times L\times d}$ (where $L$ denotes the sequence length). The encoder then combines LR Linear with lightweight deformable attention mechanisms to capture long-range dependencies:
$Z=MSDeformAttn(S,P,S)$ (8)
@@
M=LR_Linear(GELU(LR_Linear(Z)))M=LR\_Linear(GELU(LR\_Linear(Z))) (9)
@@
$Y=LR\_Linear(GELU(LR\_Linear(Y_{0})))$ (11)

### v5

# Staircase Cascaded Fusion of Lightweight Local Pattern Recognition and Long-Range Dependencies for Structural Crack Segmentation

@@ Abstract
Accurately segmenting structural cracks at the pixel level remains a major hurdle, as existing methods fail to integrate local textures with pixel dependencies, often leading to fragmented and incomplete predictions. Moreover, their high parameter counts and substantial computational demands hinder practical deployment on resource-constrained edge devices. To address these challenges, we propose CrackSCF, a Lightweight Cascaded Fusion Crack Segmentation Network designed to achieve robust crack segmentation with exceptional computational efficiency. We design a lightweight convolutional block (LRDS) to replace all standard convolutions. This approach efficiently captures local patterns while operating with a minimal computational footprint. For a holistic perception of crack structures, a lightweight Long-range Dependency Extractor (LDE) captures global dependencies. These are then intelligently unified with local patterns by our Staircase Cascaded Fusion Module (SCFM), ensuring the final segmentation maps are both seamless in continuity and rich in fine-grained detail. To comprehensively evaluate our method, this paper created the challenging TUT benchmark dataset and evaluated it alongside five other public datasets. The experimental results show that the CrackSCF method consistently outperforms the existing methods, and it demonstrates greater robustness in dealing with complex background noise. On the TUT dataset, CrackSCF achieved 0.8382 on F1 score and 0.8473 on mIoU, and it only required 4.79M parameters.
keywords:
Civil Infrastructure , Structural Crack , Lightweight Networks , Cascaded Fusion , Local Patterns , Long-Range Dependencies
journal: Engineering Structures
[…]
{graphicalabstract}[Uncaptioned image]
{highlights}
Research highlight 1
This paper proposed a novel CrackSCF network for structural crack segmentation. By introducing a lightweight convolution block, a lightweight long-range dependency extractor, and a staircase cascaded fusion module, this network effectively distinguishes between background and crack pixels with low computational resources, resulting in high-quality segmentation maps.
Research highlight 2
This paper collected the TUT dataset, which includes structural cracks in eight different scenarios, providing a more comprehensive evaluation of the network’s performance.
Research highlight 3
Comprehensive experiments conducted on five public datasets and the TUT benchmark dataset demonstrate that the proposed CrackSCF network outperforms existing SOTA methods.
@@ 1 Introduction
Cracks are common defects in structures such as pavements and buildings, which can lead to serious safety issues. Thus, automated detection on edge control devices is crucial for ensuring production safety [1, 3, 26, 18, 21]. Early automated detection methods used traditional image processing techniques for crack extraction [43, 39, 23, 31]. While these methods are easy to implement, their performance in complex scenarios still needs improvement.
@@
In recent years, benefiting from the strong local inductive capability of Convolutional Neural Networks (CNN), CNN-based networks such as MST-Net [25] and FcaHRNet [5] have shown efficient feature extraction capabilities in structural crack segmentation tasks. The spatial invariance of CNNs allows them to detect cracks anywhere in an image by learning similar features at different locations. In labeled images, crack labels are long, narrow structures of white pixels, while the background is large black areas. This challenges the model to distinguish between thin cracks and the background. Nevertheless, networks based on CNNs face challenges in capturing long-range dependencies between crack pixels, leading to segmentation discontinuities.
@@
The success of Vision Transformer (ViT) [8] and Swin Transformer [20] has demonstrated the advantages of Transformer [32], with their strong sequence processing capabilities, in handling visual tasks [42, 45]. Unlike CNNs, transformers can capture dependencies between any pixels when processing images [10, 29], making them more advantageous in handling complex topologies. To harness the advantages of both CNNs and transformers effectively, researchers have proposed networks such as CATransUNet [6], MFAFNet [7], and Crackmer [35], which integrate local and long-range information through U-shaped connections or residual connections.
Figure 1: Performance comparison of our proposed method with UCTNet [33], SFIAN [4], CT-crackseg [30], DTrCNet [35], Crackmer [34] and Simcrack [13] on the TUT dataset in terms of mIoUs, FLOPs and Params metrics.
@@
Despite the aforementioned methods combining CNN and Transformer have improved performance in pixel-level crack segmentation tasks, they still have some limitations. The different information fusion strategies significantly affect the segmentation performance of CNN-Transformer networks. For instance, UCTNet [33] and CT-crackseg [30] use channel concatenation to fuse local patterns and long-range dependences information. However, concatenation merely stacks the feature information together without deeper feature interaction, failing to understand their correlation and complementarity, resulting in poor segmentation performance on fine cracks and susceptibility to background noise. Additionally, DTrCNet [35] and Crackmer [34] use channel attention to fuse information, which can enhance focus on certain features but neglects spatial information and pixel details in the feature maps, impacting the model’s overall performance. As shown in Figure 1(a), due to these shortcomings, these four methods fail to achieve the best performance.
@@
Most networks, in pursuit of performance, overlook the significant parameter and computational costs brought by numerous convolution operations, this makes them challenging to deploy on resource-constrained control devices. As shown in Figure 1(b) and Figure 1(c), networks like SFIAN [4], Crackmer [34], and CT-crackseg[30] have very high computational FLOPs and parameter counts, indicating that they are difficult to implement on resource-limited devices like smartphones and drones, significantly diminishing their practicality.
@@
To address these issues, this paper proposes a staircase cascaded fusion network that generates high-quality segmentation maps with low computational resources while effectively suppressing background noise. Specifically, to reduce the number of parameters and computational burden, a lightweight convolution block called LRDS is introduced. This module utilizes the principle of low-rank approximation to effectively reduce the computational resources required by convolutional layers and captures spatial structures efficiently in the low-rank space through depthwise convolution, maintaining strong feature perception. The proposed architecture also includes a lightweight Long-range Dependency Extractor (LDE), which uses a lightweight deformable attention mechanism to effectively capture irregular long-range dependencies between crack pixels with minimal computational demands. Additionally, to better leverage the complementarity and correlation of feature maps while ensuring that critical detail information is not missed, the Staircase Cascaded Fusion Module (SCFM) is incorporated. This module generates feature maps layer by layer and effectively fuses them using a pixel attention mechanism that finely captures local and long-range information along with a channel concatenation mechanism, comprehensively combining local texture information and pixel dependencies.
@@
To comprehensively evaluate the structural crack segmentation performance of the model in various complex environments, this paper also collected a dataset named TUT, which contains 1408 RGB images. Existing public datasets like DeepCrack [17] and Crack500 [40] include only 1-2 types of scenarios and have relatively simple crack shapes, which are insufficient for thoroughly testing the network’s performance. The TUT dataset includes a broader range of real structural crack image scenarios, covering plastic runways, bricks, tiles, cement, bitumen, generator blades, metal materials, and underground pipelines. With such a diverse set of image scenarios, this dataset effectively evaluates the network’s generalization and robustness. Furthermore, the images feature occlusions, highly complex backgrounds, and uneven lighting, which further enhances the datasets diversity.
@@ 2.1 Network Architecture
As shown in Figure 2, the CrackSCF network has three main components: a lightweight Multi-scale Feature Extractor (MFE), a lightweight Long-range Dependency Extractor (LDE), and a Staircase Cascaded Fusion Module (SCFM). The overall architecture of the network is shaped like a staircase, with feature maps of different scales progressively fused and upsampled. In this process, the advantages of CNNs and Transformers in feature extraction complement each other.
Specifically, for a given dataset, the crack images are first input into the MFE module, which is divided into four stages. In each stage, the extracted feature maps undergo a Feature Enhancement stage. Then, the LDE module processes these enhanced feature maps to obtain pixel sequences rich in long-range dependencies. Subsequently, the SCFM module receives these sequences along with the four layers of feature maps extracted by the MFE. After processing through four stages, the resolution is doubled at each stage while the number of channels is halved, ultimately resulting in the segmentation result.
Importantly, as shown in Figure 3, using the original convolution operations results in high FLOPs and Params, especially in the SCFM module, reaching 73.36G and 9.71M, respectively. Replacing all convolutions with LRDS blocks significantly reduced FLOPs and Params, demonstrating the effectiveness of LRDS in model lightweighting.
Figure 3: Comparison of FLOPs and Params for each module using the original convolution and our proposed LRDS convolution block.
@@ 2.2 Lightweight Multi-scale Feature Extractor
As shown in Figure 4, our designed LRDS block adopts a bottleneck-like structure. It comprises a Reduce layer, a depthwise convolution layer, a pointwise convolution layer, and an Expand layer. The input feature map first passes through the Reduce layer, reducing the input channels to $C_{m}$. It then goes through the depthwise convolution layer and the pointwise convolution layer, keeping the number of channels unchanged. Finally, the Expand layer increases the number of channels to $C_{d}$, producing the final output feature map.
@@
Specifically, in the convolutional layer, the filter is assumed to have a spatial size of $k$ and a number of input channels of $c$. In computing the convolutional response, the filter is applied to a sub-tensor of the input tensor of shape $k\times k\times c$. Let this sub-tensor be denoted as $t\in\mathbb{R}^{k^{2}c+1}$, where a bias term is appended at the end. Consequently, the response at a position in the output, $y\in\mathbb{R}^{e}$, can be expressed as follows:
@@
where $e$ represents the number of filters and $W$ is an $e\times(k^{2}c+1)$ matrix. Assuming the vector $y$ lies in a low-rank subspace, it can be expressed as $y=U(y-y_{1})+y_{1}$, where $U$ is an $e\times e$ matrix of rank $e_{0}$, with $e_{0}<e$. Here, $y_{1}$ denotes the mean vector of the response. Therefore, the response can be formulated as
@@
where $b=y_{1}-Uy_{1}$ is the new bias. Given that the rank of $U$ is $e_{0}$, it can be decomposed into two $e\times e_{0}$ matrices $J$ and $K$ such that $U=JK^{\top}$. Let $W_{1}=K^{\top}W$ denote an $e_{0}\times(k^{2}c+1)$ matrix, which corresponds to a set of $e_{0}$ filters. Thus, the response can be rewritten as follows:
@@
Low-rank approximation theory significantly reduces the parameters and computational load of convolutional layers. To efficiently capture spatial information within the low-rank space for finer perception of detailed regions, the LRDS module processes features by embedding depthwise convolution and pointwise convolution directly into the low-rank space. Consequently, the MFE is constructed by replacing all standard convolution operations in the ResNet50 backbone with LRDS convolution blocks.
Figure 4: Illustration of LRDS convolution block. After the number of channels of the input features is reduced, their local and spatial features are efficiently extracted by depth-wise convolution and point-wise convolution, and finally the number of channels is restored.
@@ 2.3 Lightweight Long-range Dependency Extractor
$\centering\begin{split}&\text{MSDeformAttn}(z_{q},\hat{p}_{q},\{x_{c}\}_{c=1}^{C})=\\ &\sum_{h=1}^{H}W_{h}\Bigg[\sum_{c=1}^{C}\sum_{t=1}^{T}A_{hcqt}\cdot W^{\prime}_{h}x_{c}(\phi_{c}(\hat{p}_{q})+\Delta p_{hcqt})\Bigg]\end{split}\@add@centering$ (4)
@@
As shown in Figure 5(a), Linear operations account for 90% of the FLOPs and 67% of the Params in the LDE module. To reduce the computational resources required by the LDE, a linear transformation method (LR Linear) based on low-rank approximation is proposed. In traditional linear layers, the weight matrix connecting the input and output is typically a high-rank matrix, representing a complete connection between input and output channels. By applying low-rank approximation, the weight matrix is decomposed into two smaller matrices, effectively reducing the number of parameters.
@@
This transformation is performed through two linear layers. First, a linear transformation $B^{\top}$ is applied to the input $x\in\mathbb{R}^{r_{\text{in}}}$, resulting in an intermediate representation $z\in\mathbb{R}^{r}$:
@@
Subsequently, another linear transformation $A$ is applied to $z$, resulting in the output $y\in\mathbb{R}^{r_{\text{out}}}$:
@@
As $r\ll r_{\text{out}},r_{\text{in}}$, the low-rank approximation reduces the original matrix to two smaller matrices, significantly reducing the computational and storage requirements. This leads to a reduction in both the computational complexity and the number of parameters, thus improving training and inference efficiency. As shown in Figure 5(b), replacing all Linear operations in the LDE with LR Linear reduced the parameters and computational load from 7.18G and 1.05M to 0.14G and 0.02M.
The structure of LDE is shown in Figure 2. LDE facilitates long-range dependency modeling through the interaction of multi-scale features between the encoder and decoder. Initially, the encoder inputs a multi-scale feature set $\{X_{c}\}_{c=1}^{C}$ (where $C$ represents the number of feature levels). After undergoing layer flattening and positional encoding enhancement, the encoder extracts the feature sequence $S\in\mathbb{R}^{B\times L\times d}$ (where $L$ denotes the sequence length). The encoder then combines LR Linear with lightweight deformable attention mechanisms to capture long-range dependencies:
$Z=\text{MSDeformAttn}(S,P,S)$ (8)
@@
$M=\text{LR\_Linear}(\text{GELU}(\text{LR\_Linear}(Z)))$ (9)
@@
$Y=\text{LR\_Linear}(\text{GELU}(\text{LR\_Linear}(Y_{0})))$ (11)